Your friend works as a waitress at a local restaurant. Her income consists of an hourly wage plus tips. On Wednesday, your friend earned $25 in tips over a 5 hour period. On Friday, your friend earned $30.76 in tips over a 3 hour period. How much is your friend's hourly wage if your friend earned the same amount of money on Wednesday as on Friday?
Answer by checkley77(12844)

5X+25=3X+30.76
5X-3X=30.76-25
2X=5.76
X=5.76/2
X=2.88 IS THE HOURLY WAGE.
PROOF:
5*2.88+25=3*2.88=30.76
14.40+25=8.64+30.76
39.40=39.40